Xn
i=1
i2
i = (n − 1)2n+1 + 2
1
Expert's answer
2019-11-19T09:27:25-0500

Using induction to prove


We need prove a statement "\sum_{i=0} ^ {n} i \times 2^i = (n - 1) \times 2^{n+1} + 2"


Proof:


for n = 1


LHS =

"\sum_{i=0} ^ {1} i \times 2^i = 2"




RHS =

"(n - 1) \times 2^{n+1} + 2 = 0 + 2 = 2"


"LHS = RHS"


The given statement is true for n= 1


Let the given statement is true for n = k (integer)



"\sum_{i=0} ^ {k} i \times 2^i = (k - 1) \times 2^{k+1} + 2"




Now, Let n = k +1



"\sum_{i=0} ^ {k+1} i \times 2^i = \sum_{i=0} ^ {k} i \times 2^i + ( k+1 ) \times 2^{k+1}"

"= (k - 1) \times 2^{k+1} + 2 + (k+1 ) \times 2^ {k+1}"

"= k \times 2^ {k+1} - 2^ {k+1} + 2 + k \times 2^{k+1} + 2^{k+1}"


"= 2k \times 2^ {k+1} + 2 = k \times 2 ^ {k+1+1} + 2"

"\sum_{i=0} ^ {k+1} i \times 2^i = (k+1 -1) \times 2^ {k+1+1} + 2"

The given statement is true for n = k+1. By the principle of mathematical induction, it was proved that


"\sum_{i=0} ^ {n} i \times 2^i = (n - 1) \times 2^{n+1} + 2"